OpenStudy (2yoututu11):

What is the distance between the points (1, –6) and (–5, 2)? A. 5.6 units B. 7.2 units C. 9 units D. 10 units

OpenStudy (anonymous):

\[\LARGE d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \]

OpenStudy (anonymous):

Distance = \[\sqrt{(1+5)^{2}+(-6-2)^{2}} = \sqrt{6^{2}+8^{2}} = \sqrt{100} = 10\]